OpenStudy (anonymous):

I need a pedagogical exercise of the application of the Weierstrass \(M\)-test. Anyone?

OpenStudy (anonymous):

Here is the theorem's statement: (Weierstrass \(M\)-test) Suppose that \((f_n)\) is a sequence of functions defined on \(S\) and \(M_n\) is a sequence of nonnegative numbers such that\[|f_n(x)|\leq M_n\text{, for all }x\in S\text{ and all }n\in\mathbb{N}.\]If \(\sum M_n\) converges, then \(\sum f_n\) converges uniformly on \(S\).

OpenStudy (precal):

Sorry this is out of my league.

OpenStudy (anonymous):

Same

OpenStudy (precal):

pedagogical exercise Does that mean you are trying to teach this theorem and need a specific example?

OpenStudy (anonymous):

That's correct!

OpenStudy (anonymous):

why was i tagged in this? im retarded

OpenStudy (precal):

What type of class is this for? Just curious

OpenStudy (anonymous):

This is real analysis.

OpenStudy (precal):

Yuck Real Analysis, my nightmares are returning

OpenStudy (anonymous):

loool

OpenStudy (anonymous):

seems like firstly, if fn is <= to Mn, then it seem intuitive that if Mn converges, then fn could too. So you want an application of this idea? Like an example problem so that you could teach someone else? Like create a word problem?

OpenStudy (precal):

http://en.wikipedia.org/wiki/Uniform_convergence Can't you use something like the examples shown?

OpenStudy (anonymous):

But the only example Wikipedia gives for a uniformly convergent sequence of functions is on \(\mathbb{C}\). :(

OpenStudy (precal):

http://yourmathsolver.blogspot.com/ this looks like a cool site to check out

OpenStudy (anonymous):

Suppose\[f _{k}(x)=1/k^2(\sin(x)+\cos(x))\]So, now we have\[|f _{k}(x)|\le2/k^2=M _{k}\]Further, note\[\sum_{K=1}^{\infty}M_k<\infty\]So \[\sum_{K=1}^{\infty}|f _{k}(x)|<\infty\]

OpenStudy (anonymous):

You think that one is OK?

OpenStudy (anonymous):

Give me a second to work it out.

OpenStudy (anonymous):

That one works. :) Thanks!

OpenStudy (anonymous):

No sweat. The typing is the hardest part.....LOL